OpenStudy (anonymous):

A projectile is fired at v0 = 342.0 m/s at an angle of θ = 71.1o with respect to the horizontal. Assume that air friction will shorten the range by 32.1%. How far will the projectile travel in the horizontal direction, R?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@BAdhi

OpenStudy (gollywolly):

Calculate the horizontal and vertical components of velocity. Determine the time it will take to rise and fall due to gravity. Apply that time to horizontal speed then reduce it by the air friction.

OpenStudy (anonymous):

What formula do I need to determine that?

OpenStudy (gollywolly):

\[Cos \theta=v_x/v_0\]\[\sin \theta=v_y/v_0\]

OpenStudy (anonymous):

What would be the vx or vy?

OpenStudy (anonymous):

they would be the horizontal (\(v_x\)) and vertical (\(v_y\)) components of the initial velocity vector \(\vec{v_0}=\langle v_x, v_y\rangle \)

OpenStudy (anonymous):

So x=342.0 m/s cos(71) = 111.3443088 y=342.0 m/s sin(71) = 323.3673529

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@CShrix

OpenStudy (whpalmer4):

Isn't the angle 71.1 degrees, not 71? Your x and y components will be off a little bit if that is true.

OpenStudy (whpalmer4):

@gollywolly gave the right method, in my opinion. Find your \(v_y\) component first, so that you can write and solve the equation for the height of the projectile in order to find the time of flight. Then find the horizontal distance by multiplying the \(v_x\) component by time of flight, and reducing by the percentage mentioned for air resistance.

OpenStudy (anonymous):

x=110.78 m/s y=323.56 m/s

OpenStudy (anonymous):

@whpalmer4 I got 75.21962m

OpenStudy (anonymous):

@CShrix

OpenStudy (whpalmer4):

hmm, you're going to shoot something straight up in the air at about 700 miles per hour, how quickly do you think that will come back to the ground?

OpenStudy (anonymous):

I got 71.93s for the time.

OpenStudy (anonymous):

Scratch that. Was looking at the wrong thing. I got 65.97s

OpenStudy (whpalmer4):

if it is in the air for several dozen seconds, and it has a horizontal velocity of 110 m/s, how does it manage to only go 75 meters?!?

OpenStudy (whpalmer4):

I get 66.0327 seconds for time of flight, using g = 9.8 m/s^2

OpenStudy (anonymous):

Shouldn't it be -9.81m/s^2?

OpenStudy (whpalmer4):

the magnitude is 9.8...the time changes to 65.9653 or 65.97 if I use 9.81, that's still not explaining your horizontal error :-) If I didn't have the right sign on the first term, the two solutions for t would be 0 and -65.97 seconds...

OpenStudy (anonymous):

The way I did it was delta y =1/2(-9.81)detla t^2 +323.56 delta t 0=-4.905 delta t^2 +323.56 t divide t from t^2 and t 0=-4.905t +323.56 -323.56=-4.905t t=65.97s

OpenStudy (whpalmer4):

not sure about your variable names, but the result is fine. so the projectile is up in the air for about 66 seconds, and traveling horizontally for each of the 66 seconds at 110.78 m/s. How did you get just 75.21962m for the distance traveled (prior to accounting for air resistance)?

OpenStudy (anonymous):

0=1/2(0)(65.97)^2+110.78

OpenStudy (anonymous):

0+110.78(.321)

OpenStudy (anonymous):

110.78-35.56038

OpenStudy (anonymous):

=75.21962m

OpenStudy (anonymous):

Just follow the procedures that @whpalmer4 and @gollywolly gave!